### v1

# OSV: One Step is Enough for High-Quality Image to Video Generation

Abstract
Video diffusion models have shown great potential in generating high-quality videos, making them an increasingly popular focus. However, their inherent iterative nature leads to substantial computational and time costs. While efforts have been made to accelerate video diffusion by reducing inference steps—through techniques like consistency distillation and GAN training—these approaches often fall short in either performance or training stability. In this work, we introduce a two-stage training framework that effectively combines consistency distillation with GAN training to address these challenges. Additionally, we propose a novel video discriminator design, which eliminates the need for decoding the video latents and improves the final performance. Our model is capable of producing high-quality videos in merely one-step, with the flexibility to perform multi-step refinement for further performance enhancement. Our quantitative evaluation on the OpenWebVid-1M benchmark shows that our model significantly outperforms existing methods. Notably, our 1-step performance (FVD $171.15$) exceeds the 8-step performance of the consistency distillation based method, AnimateLCM (FVD $184.79$), and approaches the 25-step performance of advanced Stable Video Diffusion (FVD $156.94$).
Introduction
Video synthesis provides rich visual effects and creative expression for films, television, advertisements, and games. Diffusion models are playing an increasingly important role in video synthesis (Blattmann et al. 2023c; Shi et al. 2024; Blattmann et al. 2023a; Singer et al. 2022; Ho et al. 2022). Typically, diffusion models involve a forward process and a reverse process. In the forward process, real data is iteratively perturbed with noise until it converges to a simple noise distribution, typically Gaussian. In the reverse process, the noise is gradually removed, ultimately transitioning back to the target data distribution. However, this reverse process usually requires the numerical solution of a generative ODE, termed Probability Flow ODE (PF-ODE) (Song et al. 2020). The iterative nature of this numerical solving process leads to significantly higher computational costs compared to other generative models (e.g., GANs) (Goodfellow et al. 2014). These computational demands are even more significant in video synthesis; for instance, generating a short 2-second video clip using Stable Video Diffusion (SVD) (Blattmann et al. 2023a) on a high-performance A100 GPU can take over 30 seconds.
Figure 1: OSV is a two-stage video diffusion acceleration strategy. In the first stage, we apply GAN for better training efficiency. In the second stage, we apply consistency distillation to boost the performance upper-bound.
Figure 2: Summative motivation. We observe and summarize crucial limitations for (latent) consistency models, and generalize the design space, well tackling all these limitations.
Currently, several strategies have been proposed to reduce the computational costs associated with video generation. Some approaches utilize consistency distillation in the latent space (LCM) (Luo et al. 2023; Song et al. 2023) for acceleration; however, they often struggle to achieve competitive results in few-step settings, such as one or two steps. Other methods initialize with pre-trained diffusion models and incorporate adversarial loss in a GAN-based framework to accelerate generation (GAN). Nonetheless, these methods frequently suffer from training instability. Moreover, since advanced diffusion models are typically trained in the latent space (i.e., they apply an auto-encoder to encode the high-resolution videos into the latent space to facilitate the training), GAN-based methods require decoding latent features into the real image or video space before passing them to the discriminator. This process incurs substantial memory usage and computational overhead, particularly in high-resolution generation tasks. Figure 2 illustrates our summative observations of previous methods, which we summarize as follows:
Exposure Issues: Although LCM can accept classifier-free guidance (CFG) (Ho and Salimans 2021), higher CFG values can lead to exposure problems (Wang et al. 2024a). This complicates the selection of hyperparameters for training.
Efficiency: LCM faces slow training convergence and often produces poor results, particularly when with less than four inference steps, which limits sampling efficiency. While introducing GANs can alleviate some of the efficiency issues of LCM, it introduces new problems.
Increased Training Burden: Because advanced diffusion models are typically trained in the latent space—where high-resolution videos are encoded via an auto-encoder—GAN-based approaches need to decode these latent features back into the real image or video space before they can be evaluated by the discriminator (Sauer et al. 2023b). This decoding step significantly increases memory consumption and computational load, particularly when dealing with high-resolution or long video-length generation tasks.
Training Instability: GANs are known for training instability. Introducing GANs can easily lead to training instability, causing the model to degrade as training progresses.
Inability to Iterative Refine: GANs training does not allow for iterative refinement of generated results, unlike consistency models, leading to poorer results with more sampling iterations.
In this work, we present OSV (One Step Video Generation), allowing for high-quality image-to-video generation in one step while still supporting multi-step refinement. OSV is a two-stage video diffusion acceleration training strategy. Specifically, in the first stage, we incorporate Low-Rank Adaptation (LoRA) (Hu et al. 2021) and fully utilize GANs training, with real data as the true condition for the GANs, greatly accelerating model convergence. In the second stage, we introduce the LCM training function, with data generated by the teacher model serving as the true condition for the GANs, while only fine-tuning specific layers of the network. To further facilitate the training convergence, we replace the commonly applied one-step ODE solver with multi-step solving, ensuring higher distillation accuracy. The second training stage addresses GANs training instability in the later training phase and further improves the performance upper-bound with the knowledge transferred from the teacher video diffusion models. To further improve the training efficiency, we revise the current popular discriminator designs and propose to discard the VAE decoder in the adversarial training. As illustrated in Figure 3, we replace the VAE decoder with a simple up-sampling operator. That is, we directly feed the upsampled video latent into the discriminator whose backbone is pre-trained on the real image/video space (DINOv2 (Oquab et al. 2023)). Our ablation study shows it not only reduces the training cost but also achieves better performance. Our proposed two-stage training not only facilitates early stage training efficiency but also stablizes training and improves the performance upper-bound in the later stage. As shown in Figure 1 verifies our claim.
In summary, we investigate the limitations of previous consistency model-based and GAN training-based methods and propose a two-stage training approach, combining the strength of both and achieving state-of-the-art performance on fast image-to-video generation.
Related Works
Diffusion Distillation. Denoising process usually has many steps, making them 2-3 orders of magnitude slower than other generative models such as GANs and VAEs. Recent progress on diffusion acceleration has focused on speeding up iteratively time-consuming generation process through distillation (Song et al. 2023; Song and Dhariwal 2023; Salimans and Ho 2022; Heek, Hoogeboom, and Salimans 2024; Yan et al. 2024; Ren et al. 2024; Zheng et al. 2024; Sauer et al. 2023b, 2024; Kim et al. 2023; Zhou et al. 2024; Wang et al. 2024a). Typically, they train a generator to approximate the ordinary differential equation (ODE) sampling trajectory of the teacher model, resulting in fewer sampling steps. Particularly, Progressive Distillation (Salimans and Ho 2022; Meng et al. 2023) trains the student to predict directions pointing to the next flow locations. ADD (Sauer et al. 2023b) leverages an adversarial loss to ensure high-fidelity image generation. SDXL-Lighting (Lin, Wang, and Yang 2024) combines progressive progressive distillation and adversarial distillation, striking a balance between mode coverage and quality.
When it comes to video distillation, Video-LCM (Wang et al. 2023) builds upon existing latent video diffusion and incorporates consistency distillation techniques, achieving high-fidelity and smooth video synthesis with only four sampling steps. Furthermore, Animate-LCM (Wang et al. 2024b) proposes a decoupled consistency learning strategy that separates the distillation of image generation priors and motion generation priors, enhancing visual quality and training efficiency. SF-V (Zhang et al. 2024) follows diffusion-as-GAN paradigms and proposes single-step video generation by leveraging adversarial training on the SVD (Blattmann et al. 2023b) model. In contrast, our empirical observations show that decoupling adversarial latent training and consistency distillation improves training stability and generation quality. Additionally, instead of using UNet itself as the discriminator feature encoder, we use DINOv2 (Oquab et al. 2023) as a feature extractor, which both improves training efficiency and generation quality. Moreover, a multi-step consistency distillation training strategy is also proposed.
Preliminaries
Diffusion Models. Diffusion mdoel (Ho, Jain, and Abbeel 2020; Song et al. 2020) gradually introduces random noise through the diffusion process, transforming the current state ${\mathbf{x}_{0}}$ into a previous state ${\mathbf{x}_{t}}$.
We consider the continuous case of diffusion models. The forward process of a diffusion model can be described as:
$d{\mathbf{x}}={f}_{t}({\mathbf{x}})\,dt+g_{t}\,d\boldsymbol{w}$ (1)
where $f_{t}(\mathbf{x})=\frac{d\log\alpha_{t}}{d{t}}\mathbf{x}$ and $g_{t}^{2}=\frac{\mathrm{d}\sigma_{t}^{2}}{dt}-2\frac{d\log\alpha_{t}}{dt}\sigma_{t}^{2}$. ${\alpha}_{t}$ is the predefined scale factor and $\boldsymbol{w}_{t}$ denoting the standard Winer process. $\sigma_{t}$ controls the level of noise.
Considering the reverse process of diffusion models in the continuous case:
$d\mathbf{x}=\left({f}_{t}(\mathbf{x})-\frac{1}{2}g_{t}^{2}\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})\right)\,dt$ (2)
This is known as the Probability Flow Ordinary Differential Equation (PF-ODE). We use a neural network $\epsilon_{\theta}(\mathbf{x}_{t},t)$ to approximate $\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})$.
Consistency Models. Consistency Models (Song et al. 2023) are built upon the PF-ODE in continuous-time diffusion models. Given a PF-ODE that smoothly transforms data into noise, Consistency Models learn to map any point to the initial point of the trajectory at any time step for generative modeling. The formula can be described as:
${f}:\left(\mathbf{x}_{t},t\right)\mapsto\mathbf{x}_{\epsilon},\quad t\in[\epsilon,T].$ (3)
To ensure the boundary conditions hold for any consistent functions, Consistency Models typically employ skip connections. Suppose we have a free-form deep neural network $F_{\theta}$, which can be formulated as:
$f_{\theta}(\mathbf{x},t)=c_{\text{skip}}(t)\mathbf{x}+c_{\text{out}}(t)F_{\theta}(\mathbf{x},t)$ (4)
where $c_{\text{skip}}(t)$ and $c_{\text{out}}(t)$ are differentiable functions such that $c_{\text{skip}}(\epsilon)=1$, and $c_{\text{out}}(\epsilon)=0$.
For training Consistency Models, the output is enforced to be the same for any pair belonging to the same PF-ODE trajectory, i.e., $f(\mathbf{x}_{t},t)=f(\mathbf{x}_{t^{\prime}},t^{\prime})$ for all $t,t^{\prime}\in[\epsilon,T]$. To maintain training stability, an Exponential Moving Average (EMA) of the target model is used, given by:
$\mathcal{L}_{\text{CD}}^{N}(\theta,\theta^{-};\phi)=\mathbb{E}\left[\lambda(t_{n})d\left(f_{\theta}(\mathbf{x}_{t_{n+1}},t_{n+1}),f_{\theta^{-}}({\mathbf{x}}_{t_{n}}^{\phi},t_{n})\right)\right]$
where $\lambda(t_{n})$ is a weighting function, $d(\cdot,\cdot)$ is a distance metric and ${\mathbf{x}}_{t_{n}}^{\phi}=\mathbf{x}_{t_{n+1}}+(t_{n}-t_{n+1})\Phi(\mathbf{x}_{t_{n+1}},t_{n+1};\phi)$. $\Phi(\cdots;\phi)$ represents the update function of a one-step ODE solver applied to the empirical PF-ODE. $\theta^{-}$ is the EMA version weight of target models.
Figure 3: Comparison of Different Adversarial Training Methods. SF-V requires the encoder of UNet as the feature extraction part of the discriminator. ADD perform adversarial distillation on raw image pixel, which needs to convert latent to image thorough VAE Decoder. Differently, we directly up-sampling the latents, thus requiring less memory and computational resources due to the absence of VAE Decoder.
Figure 4: Overview of OSV. In the first stage, we combine GAN loss and Huber loss (Song and Dhariwal 2023) for better training efficiency. In the second stage, we use consistency distillation loss to boost the performance upper-bound. $b_{sahpe}$, $t_{shape}$, $c_{shape}$, $h_{shape}$ and $w_{shape}$ represent the batch size, number of frames, color channels, height, and width of the input video, respectively.
Method
In this section, we introduce the specific technical details of our OSV model. The model employs a two-stage training method to minimize GAN training instability and incorporates a novel multi-step consistency model solver to enhance its efficiency. Observing the negative impact of CFG on the distilled model, we remove CFG and design a new high-order solver.
In this section, we detail our OSV model, which employs a two-stage training process to enhance video generation. The first stage leverages GAN training, allowing for rapid improvement in image quality during the initial training steps. The second stage combines GAN training with consistency distillation, providing a balanced approach that further stabilizes training and enhances model performance. Finally, we introduce a novel high-order solver, which refines generation results by a high-order prediction, leading to higher accuracy and efficiency in video generation.
Network Components. The OSV training process consists of three main components: a student model with weights $\theta$, a EMA model with weights $\theta^{-}$, a pre-trained teacher model with frozen weights $\phi$, and a discriminator with weights $\psi$, as shown in Figure 4. Specifically, the student and teacher models share the same architecture, with the student model initialized from the teacher model. For the discriminator, we adopt the same structure as StyleGAN-T (Sauer et al. 2023a), utilizing DINOv2 (Oquab et al. 2023). We freeze the pretrained weight of DINOv2 and add trainable temporal discriminator heads and spatial discriminator heads for discrimination of the features extracted DINOv2 inspired by SF-V (Zhang et al. 2024). The temporal discriminator heads are composed of 1D convolution blocks. The spatial discriminators are composed of 2D convolution blocks.
Latent GAN Pretraining. As shown in Figure 3, SF-V and ADD implement different adversarial distillation methods. The discriminator in SF-V shares the same architecture and weights as the pre-trained UNet encoder backbone and is enhanced by adding a spatial discrimination head and a temporal discrimination head after each backbone block. ADD uses DINOv2 as the discriminator. Although the discriminator in ADD reduces computational load and memory usage compared to SF-V, the student model’s generated data needs to be passed through a VAE decoder before being input into the discriminator, which undoubtedly increases both computational load and memory usage. We find that using DINOv2 directly in the latent space also achieves adversarial distillation and significantly saves memory and computational resources compared to the pixel space.
We find that during the pre-training phase of the network, introducing only GAN for adversarial distillation, without LCM, can achieve rapid image quality improvement at low steps. Specifically, the student model optimizes the Huber loss and adversarial loss between the generated data $f_{\theta}(\mathbf{x}_{t_{n}},t_{n})$ and the real data $\mathbf{x}_{0}$, as follows:
$\mathcal{L}_{\text{OSV}}^{\text{d}_{1}}(\boldsymbol{\theta},\boldsymbol{\theta}^{-};\boldsymbol{\phi};\boldsymbol{\psi})$ =$\text{ReLU}(1-D_{\psi}({\mathbf{x}}_{0}))+\text{ReLU}(1+D_{\psi}(f_{\theta}(\mathbf{x}_{t_{n}},t_{n})))$
$\mathcal{L}_{\text{OSV}}^{\text{g}_{1}}(\boldsymbol{\theta},\boldsymbol{\theta}^{-};\boldsymbol{\phi};\boldsymbol{\psi})$ = $\lambda^{LGP}*\text{ReLU}(1-D_{\psi}(f_{\theta}(\mathbf{x}_{t_{n}},t_{n})))+d(\mathbf{x}_{0},f_{\theta}(\mathbf{x}_{t_{n}}^{\phi},t_{n}))$
where $D_{\psi}$ is the discriminator, $\text{ReLU}(x)=x$ if $x>0$ else $\text{ReLU}(x)=0$ and $\lambda^{LGP}$ is a hyper-parameter. $d(x,y)=\sqrt{\parallel x-y\parallel^{2}_{2}+c^{2}}-c$ (Song and Dhariwal 2023), where $c>0$ is an adjustable constant. It is also worth noting that during this stage, we load the student model with LoRA training and train for a very short period. This ensures rapid image quality improvement at low steps without causing training instability in the student model. Given the mismatch between the pre-training data distribution and the distillation data distribution, using LoRA ensures that the student model retains most of the teacher model’s knowledge while reducing the degradation of image quality at low steps.
Adversarial Consistency Latent Distillation. In the second stage, we introduce the consistency loss. We initialize the student model and EMA model for this stage using the student model from the first stage. Similarly, we initialize the discriminator for the second stage using the discriminator from the first stage. Unlike AnimateLCM, we solve for $\mathbf{x}_{t_{n}}^{\phi}$ using the following equations:
[…]
where $c$ represents the condition, which in our case is an image embedding. $m$ controls the number of steps of the ODE solver. We introduce classifier-free guidance distillation similar to (Luo et al. 2023): $\hat{\Phi}(\mathbf{x}_{t_{n+m}},t_{n+m},c;\phi)={\Phi}(\mathbf{x}_{t_{n+m}},t_{n+m},c_{zero};\phi)+w*({\Phi}(\mathbf{x}_{t_{n+m}},t_{n+m},c;\phi)-{\Phi}(\mathbf{x}_{t_{n+m}},t_{n+m},c_{zero};\phi))$, where $c_{zero}$ represents the image embedding set to zero and $w$ controls the strength. We also optimize the adversarial loss between the student-generated data $f_{\theta}(\mathbf{x}_{t_{n+m}})$ and the teacher-generated data $f_{\theta^{-}}(\mathbf{x}_{t_{n}}^{\phi},t_{n})$, as follows:
$\mathcal{L}_{\text{OSV}}^{\text{d}_{2}}(\boldsymbol{\theta},\boldsymbol{\theta}^{-};\boldsymbol{\phi};\boldsymbol{\psi})$ =$\text{ReLU}(1-D_{\psi}(f_{\theta^{-}}(\mathbf{x}_{t_{n}}^{\phi},t_{n})))+\text{ReLU}(1+D_{\psi}(f_{\theta}(\mathbf{x}_{t_{n+m}},t_{n+m})))\,,$ (9)
$\mathcal{L}_{\text{OSV}}^{\text{g}_{2}}(\boldsymbol{\theta},\boldsymbol{\theta}^{-};\boldsymbol{\phi};\boldsymbol{\psi})$ = $\lambda^{ACD}*\text{ReLU}(1-D_{\psi}(f_{\theta}(\mathbf{x}_{t_{n+m}},t_{n+m})))+\lambda(t_{n})d(f_{\theta^{-}}(\mathbf{x}_{t_{n}}^{\phi},t_{n}),f_{\theta}(\mathbf{x}_{t_{n+m}},t_{n+m}))\,.$ (10)
where $\lambda^{ACD}$ is a hyper-parameter, $d(x,y)=\sqrt{\parallel x-y\parallel^{2}_{2}+c^{2}}-c,c>0$ and $\lambda(t_{n})$ is a weighting function, similar to (Song and Dhariwal 2023). It is worth noting that, as shown in Figure 5, our multi-step solving method not only achieves higher accuracy than the single-step solving method with the same number of iterations but also demonstrates higher accuracy within the same training time. (When $m$ is set to 5, the time for 22,000 iterations is approximately equal to the time for 42,000 iterations when $m$ is set to 1.)
Figure 5: Effectiveness of the proposed Time Travel Sampler. Compared to one solver step, multi-step solving exhibits a faster training convergence speed and superior performance, demonstrating the effectiveness of the proposed method. $m$ is set to 5.
High Order Sampler Based On Time Travel. As shown in Figure 2, we observed the negative impact of using CFG on the distilled model. Even with smaller CFG weights, the improvement in video generation was minimal. We decided to remove CFG and design a higher-order sampler named Time Travel Sampler (TTS). Suppose the number of sampler steps is set to $k$, corresponding to the time function $t^{k}$, and the number of sampler steps is set to $k+1$, corresponding to the time function $t^{k+1}$. $t^{k+1}$ has one more time step compared to $t^{k}$. Let $t^{k}_{0}=0$ and $t^{k+1}_{0}=0$, we have $t^{k+1}_{i}<t^{k}_{i},t^{k+1}_{i}>t^{k}_{i-1},\forall i\in[1,k]$. During sampling, we set the number of sampler steps to $k$. Observing that the image generation quality improves as the time step $t$ decreases, at step $i$, given the sample $x_{t^{k}_{i}}$, we can first solve for the sampling result at the lower time step $t^{k+1}_{i}$ and then revert to $t^{k}_{i}$ to solve the consistency function again:
$&f_{\theta}(\mathbf{x}_{t^{k}_{i}},t^{k+1}_{i})=c_{\text{skip}}(t^{k+1}_{i})\mathbf{x}_{t_{i}^{k}}+c_{\text{out}}(t^{k+1}_{i})F_{\theta}(\mathbf{x}_{t^{k}_{i}},t^{k+1}_{i}),\\ &\hat{\mathbf{x}}_{t^{k}_{i}}=c_{\text{skip}}(t^{k}_{i})f_{\theta}(\mathbf{x}_{t^{k}_{i}},t^{k+1}_{i})+c_{\text{out}}(t^{k}_{i})F_{\theta}(f_{\theta}(\mathbf{x}_{t^{k}_{i}},t^{k+1}_{i}),t^{k}_{i}).\end{aligned}$
Training Data Resolution. We found that setting different training data sizes for different training stages can accelerate the training process. In the first stage, the quality of the generated videos is evidently related to the dataset size, so we use a resolution of 1024$\times$576. In the second stage, we train the model using a resolution of 576$\times$320.
Experiment
Implementation details. We apply stable video diffusion as the base model for most experiments. we uniformly sample 100 timesteps for training. For better training and evaluation of our method, we utilize OpenWebVid-1M (Nan et al. 2024) for training and validation. We randomly select 1 million videos from OpenWebVid-1M as the training set and 1,000 videos as the test set. In the first stage, we fixed the resolution of the training videos at 1024$\times$768, the FPS at 7, the batch size at 1, and the learning rate at 5e-6. The training was conducted over 2K iterations on 8 NVIDIA H800 GPUs. In the second stage, we fixed the resolution of the training videos at 576$\times$320, the FPS at 7, the batch size at 1, and the learning rate at 5e-6. The training was conducted over 20K iterations on 2 NVIDIA H800 GPUs. All stages use the Adam optimizer (Kingma and Ba 2014). We use Fréchet Gesture Distance (FVD) (Unterthiner et al. 2018) to evaluate our model.
Figure 6: Qualitative generation results. One-step result of OSV with TTS achieves superior video clarity compared to other baselines.
Table 1: Image-to-video performance comparison on the validation set of OpenWebVid-1M. $\dagger$ means our implementations of SF-V (no public weight).
Name Steps $\downarrow$ NFE $\downarrow$ FVD $\downarrow$
[…]
8 16 229.94
2 4 1015.25
1 2 1177.34
AnimateLCM (Wang et al. 2024b) 8 8 184.79
4 4 405.80
2 2 575.33
1 1 997.94
SF-V (Zhang et al. 2024$\dagger$ 1 1 425.68
Ours 1 1 335.36
[…]
Ours 2 2 181.95
Table 2: Ablation study with OSV. Unless otherwise specified, we set the sampling steps to 1 and use the Time Travel Sampler(TTS).
[…]
$1$ 332.25
$5$ 171.15
(a) Effect of Multi-Step Solving.
[…]
221.75
171.15
(b) Effect of First Stage Training.
[…]
      ✗       ✗       298.35
      ✗       ✓       234.13
      ✓       ✓       171.15
(c) Effect of Second Stage Training.
[…]
405.41
171.15
(d) Effect of Adversarial Training.
[…]
      ✗       232.25
      ✓       171.15
(e) Effect of VAE Decoder.
[…]
      3.0       531.23
      1.5       426.71
      No CFG       335.36
(f) Effect of CFG.
[…]
      576$\times$320       455.35
      1024$\times$576       388.86
(g) Effect of Data Resolution in Stage One.
[…]
      576$\times$320       171.15
      1024$\times$576       181.95
(h) Effect of Data Resolution in Stage Two.
Quantitative Experiments
Table 6 illustrates the quantitative comparison of OSV with the strong baseline methods Animate-LCM and SF-V. Observing the negative impact of CFG on the distilled models, we remove CFG from both the Animate-LCM and OSV models. This reduction led to decreased inference time for the student network. OSV significantly outperforms the baseline methods, especially at low steps. Our method is compared with SF-V and AnimateLCM, using different numbers of sampling steps for each method. Additionally, we found that using the high-order sampler in a single step resulted in an FVD of 171.15, compared to an FVD of 181.95 when performing direct inference with two steps. This demonstrates better results than direct two-step inference.
Figure 7 shows the results of our user study, where our model achieved higher clarity and smoothness compared to the teacher.
Qualitative Results
Figure 6 shows the generated results of our method in image-to-video generation, all of which achieved satisfactory performance. The generated results demonstrate that our method effectively adheres to the consistency properties across different inference steps, maintaining similar styles and motions. Other methods either suffer from overexposure issues or blurring due to motion. We exhibit good visual quality and smooth motion with only one step.
Figure 7: User study comparing our distilled model with its teacher and competing distillation baselines. Animate-LCM uses 8 sampling steps, SF-V uses 1 sampling step, OSV uses 1 sampling step with TTS, Teacher uses 25 sampling steps. Supplementary materials provide more generated videos.
Ablation Studies
Effect of Multi-Step Solving. We set the OSV model with 5 solver steps as Baseline-1. To verify the effectiveness of the multi-step solving method, we remove the multi-step solving component and train the OSV model with 1 solver step under the same training settings. As shown in Table 2a, the multi-step solving method achieves higher solving accuracy. Considering the training time for all models is the same, our method also performs better, as illustrated in Figure 5.
Effect of First Stage of Training and Second Stage of Traning. We set the OSV model trained with both Stage 1 and Stage 2 as Baseline-2. To verify the effectiveness of Stage 1 training, we removed Stage 1 training and trained the OSV model with only Stage 2 under the same training settings. As shown in Table 2b, it is evident that performing both Stage 1 and Stage 2 training contributes more to the model’s convergence. Stage 1 training primarily ensures that the student model can still generate detailed content at low steps, which aids in the consistency training of Stage 2 and prevents blurring issues similar to those in Animate-LCM. To verify the effectiveness of the second stage of training, we remove the second stage and train the OSV model with only the first stage under the same training settings. As shown in Table 2c, it is evident that the second stage of training contributes more to refining the generated videos. The consistency training in the second stage further enhances the details of the generated videos. We also find that our TTS sampler is effective for distilling the student model using GAN methods.
Effect of Adversarial Training. We set the OSV model with adversarial loss as Baseline-3. To verify the effectiveness of adversarial distillation, we remove the adversarial loss and train the OSV model with only Huber loss and consistency loss under the same training settings. As shown in Table 2d, adversarial training results in higher generation quality. Using only consistency loss leads to a fitting error between the student model and the teacher model.
Effect of VAE Decoder. We set the OSV model trained with only the first stage as Baseline-4. We add the VAE Decoder from the ADD method, as shown in Figure 3, to the discriminator and trained the OSV model with only the first stage. As shown in Table 2e, we find that adding the VAE Decoder resulted in even worse performance. This indicates that performing adversarial training in the latent space is more beneficial for the discriminator. After removing the Vae Decoder, the training speed on the NVIDIA H800 improved from an average of 4.29 seconds/iteration to 2.61 seconds/iteration. In the supplementary materials, we visualize and analyze the differences between latent space videos and pixel space videos.
Effect of CFG. As shown in Table 2f, we investigate the impact of CFG on video generation by the model. Figure 2 illustrates the overexposure issue caused by adding CFG. It is evident that even reducing the CFG scale still negatively affects the model. Removing CFG not only saves time but also improves the quality of the generated videos.
Effect of Data Resolution in Stage One and Stage Two. As shown in Tables 2g and  2h, we find that different training data Resolution can be used in the first and second stages. For the first stage, we did not introduce the consistency distillation loss, so the quality of the generated videos is related to the dataset size. If we use a dataset size of 576×320, the videos are downsampled quite small, resulting in significant information loss. In the second stage, we enforce the consistency of the student model’s trajectories at different time steps, so the quality of the generated videos depends on the quality of the teacher model’s videos, and a larger video resolution is not required. Additionally, since the computational load in the first stage is much lower than in the second, we use a dataset size of 1024$\times$576 in the first stage and 576$\times$320 in the second. This significantly reduces the model’s distillation time.
Conclusion
In this paper, we introduce the OSV, which utilizes a two-stage training process to enhance the stability and efficiency of video generation acceleration. In the first stage, we employ GAN training, achieving rapid improvements in generation quality at low steps. We propose a novel video discriminator design where we leverage pretrained image backbones (DINOv2) and lightweight trainable temporal discriminator heads and spatial discriminator heads. We also innovatively propose to replace the commonly applied VAE decoding process with a simple up-sampling operation, which greatly facilitates training efficiency and improves model performance. The second stage integrates consistency distillation, further refining the model’s performance and ensuring training stability. Additionally, we show that applying multi-step ODE solver can increase the accuracy of predictions but also facilitate faster training convergence. By removing the CFG and introducing the Time Travel Sampler (TTS), we are able to further improve video generation quality. Our experiments demonstrate that the OSV significantly outperforms existing methods in both speed and accuracy, making it a robust and efficient solution for video generation acceleration.
Limitations. We observe that the distilled model exhibits less motion and heavily relies on the input images. This is a common phenomenon in distilled models with few-step sampling.